# Covariate balancing estimation and model selection for difference-in-differences approach

### v1

@@ Abstract
In causal inference, remarkable progress has been made in difference-in-differences (DID) approaches to estimate the average effect of treatment on the treated (ATT). Of these, the semiparametric DID (SDID) approach incorporates a propensity score analysis into the DID setup. Supposing that the ATT is a function of covariates, we estimate it by weighting the inverse of the propensity score. As one method to make the estimation robust to the propensity score modeling, we incorporate covariate balancing. Then, by attentively constructing the moment conditions used in the covariate balancing, we show that the proposed estimator has doubly robustness. In addition to the estimation, model selection is also addressed. In practice, covariate selection is an essential task in statistical analysis, but even in the basic setting of the SDID approach, there are no reasonable information criteria. Therefore, we derive a model selection criterion as an asymptotically bias-corrected estimator of risk based on the loss function used in the SDID estimation. As a result, we show that a penalty term is derived that is considerably different from almost twice the number of parameters that often appears in AIC-type information criteria. Numerical experiments show that the proposed method estimates the ATT robustly compared to the method using propensity scores given by the maximum likelihood estimation (MLE), and that the proposed criterion clearly reduces the risk targeted in the SDID approach compared to the intuitive generalization of the existing information criterion. In addition, real data analysis confirms that there is a large difference between the results of the proposed method and the existing method.
Keywords: Covariate balancing, Difference-in-differences, Information criterion, Model selection, Propensity score, Semiparametric approach, Statistical asymptotic theory
@@ 1 Introduction
The difference-in-differences (DID) approach is one of the most common methods used in fields such as economics and epidemiology to assess the effect of a particular intervention or treatment. In the basic setting, there are two groups: a treatment group receiving a specific treatment and a control group receiving no treatment or other treatment. The pre and post-treatment outcomes of each group are measured and the differences in the outcomes are compared between the groups to estimate the average treatment effect on the treated (ATT). Under the parallel trend assumption, simple estimation is valid. More recently, Abadie (2005) dealt with inference under more naturalized assumptions using information from the covariates, specifically the parallel trend assumption conditioned on the covariates. Then, as a method that does not necessarily require modeling of the covariates to the outcome, they propose the semiparametric DID (SDID) method, which estimates ATT by weighting the inverse of the propensity score.
@@
With or without conditionality, the parallel trend assumption is a key assumption in the DID. This is the assumption that the mean outcomes of a treatment group and a control group would have been similar over time if there had been no treatment. On the other hand, especially in epidemiology, the ignorable treatment assignment assumption has usually been invoked in causal inference (e.g., Robins et al. 1994 and Li et al. 2018). This is the assumption that the potential outcome variable and the treatment assignment variable, which represents the received treatment, become independent conditioned on the covariates. Although neither assumption is verifiable, the validity of the parallel trend assumption can be examined by visual inspection when we have observations over many time points. And the parallel trend assumption naturally holds if it can be assumed that there is no change over time in the expected value of the outcome without the treatment. Needless to say, the parallel trend assumption is considerably weaker than the ignorable treatment assignment assumption, and estimating the ATT under such desirable conditions is the advantage that the DID gained in exchange for collecting pre- and post-treatment outcomes.
@@
The methodological development of the DID has been remarkable in recent years because of such advantages and the increasing amount of available observational study data (see Roth et al. 2023 for a comprehensive review of the DID in recent years). However, for example, Abadie (2005) assumes that the model for the propensity score is correctly specified, but if the model is incorrect, the resulting estimator has a bias. As an extension of Abadie (2005), Sant’Anna and Zhao (2020) assumes not only a regression model for the treatment assignment variable used to obtain the propensity score, but also a regression model for the outcome, and proposes the estimator for the ATT that combines the results of each estimator. Their proposed estimator is consistent if either a propensity score or outcome regression working model is correctly specified, that is, the estimators have doubly robustness.
@@
In this article, as with Abadie (2005), we assume that ATT is a function of covariates and address the semiparametric estimation problem for that function. We assume two regression models as above and give double robustness to the estimation, but unlike the Sant’Anna and Zhao (2020), we incorporate covariate balancing (Imai and Ratkovic 2014), a method that uses propensity scores such that a certain moment condition is satisfied. This method actually estimates the model for the treatment assignment, while assuming, but not estimating, the model for the outcome. Nevertheless, the SDID estimator derived by weighting the inverse of the propensity score, which is obtained by the covariate balancing, will be consistent if either one of the regression models is correctly specified. One point to note is that, in order to satisfy this favorable asymptotic property in DID, the moment condition is constructed in a different way from the usual one. Incidentally, incorporating covariate balancing leads to the estimation of causal effects via weights that always balance the moments of the covariates across groups, which is also preferable from an interpretive point of view (Imai and Ratkovic 2014). We call the proposed methodology the covariate balancing for DID (CBD) method.
@@
Obviously, model selection is an essential task in the DID as well, in a series of data analysis tasks. For example, in estimating the effect of union membership on wages, Houngbedji (2016) treats a linear model with workers’ demographics for causal effects, and explores factors affecting the causal effect on the basis of the values of the regression coefficients and its $p$-value. In other words, the issue of covariate selection is addressed. However, for the information criterion that is the basic tool for model selection, there are no reasonable ones, even for the basic setting of the SDID. Thus, we derive the information criterion for the basic SDID and CBD methods as an asymptotically unbiased estimator of the risk based on the loss function used in the SDID estimation. Note that traditional methods such as GIC (Konishi and Kitagawa 1996) cannot be applied because the loss function includes weights based on the propensity score, which is a random variable. Since the SDID method places importance on being based on the conditional parallel trend assumption, it does not rely on the ignorable treatment assignment condition when deriving the information criterion. This causes some difficulties. As a result, we derive penalty terms that are significantly different from twice the number of parameters, which often appear in AIC-type information criteria including GIC.
@@
In summary, the aims of this article are to develop a new doubly robust estimation method with the advantage of covariate balancing in the SDID method, and to derive model selection criteria for the SDID method, for which no reasonable model selection criteria exist yet. The organization of this article is as follows. First, in section 2, the model and estimation target are specified, and the SDID method proposed in Abadie (2005) is introduced. In section 3, we propose the CBD method and give a theorem showing its doubly robustness. Numerical experiments show that the estimation is robust against model misspecification. In section 4, when the propensity score is known, we derive the model selection criterion as a bias-corrected estimator of the risk that is naturally defined from the loss function used in the SDID method. Then, we also give a theorem that suggests that the penalty term of the model selection criterion is asymptotically valid. As a comparison, we derive the model selection criterion the QICw, which naturally extends the idea of Platt et al. (2013) to the SDID method. Furthermore, through numerical experiments, we verify the approximation accuracy of the penalty term of the proposed criterion from the perspective of estimating the risk, and also verify its performance compared to the QICw from the viewpoint of risk minimization. Section 5 summarizes the results for the CBD method in the same manner as Section 4. In other words, model selection criteria are derived, theorems are given to support their validity, and their performance is evaluated through numerical experiments. Section 6 reports the results of estimating causal effects and model selection for the real data in LaLonde (1986). Finally, Section 7 discusses the future development of the proposed method.
@@ 2.1 Model and Assumptions
Suppose that there are two groups, a treatment group and a control group, and that outcomes are collected before and after the possible treatment time point, respectively. We treat the counterfactual setting, where $y^{[1]}(t)$ denotes the outcome of the treatment group, $y^{[0]}(t)$ the outcome of the control group, $t=0$ the outcome before treatment, and $t=1$ the outcome after treatment. In practice, only one of the outcomes of the treatment or control group can be obtained. Preparing treatment assignment variables such that $d^{[1]}=1\ (d^{[0]}=0)$ if assigned to the treatment group and $d^{[0]}=1\ (d^{[1]}=0)$ if assigned to the control group, the observed outcome collected at each time point can be expressed as
@@
Letting the change in the outcome variable at two time points in the treatment group be $\Delta^{[1]}=y^{[1]}(1)-y^{[1]}(0)$ and the change in the control group be $\Delta^{[0]}=y^{[0]}(1)-y^{[0]}(0)$, the observed change can be expressed as
@@
we suppose $g(\bm{x};\bm{\theta})$ $(\bm{\theta}\in\bm{\Theta}\subset\mathbb{R}^{p})$ as a working model to approximate it, and for simplicity, we suppose the working model to be a linear model $\bm{x}^{{\rm T}}\bm{\theta}$. Then, we estimate $\bm{\theta}$ and give the ATT by taking the expectation with respect to $\bm{x}$ in the treatment group.
Now, we give the optimal value of $\bm{\theta}$ as
@@
Abadie (2005) shows that $\bm{\theta}^{*}$ can be consistently estimated by the SDID method using propensity scores $e^{[1]}(\bm{x})\equiv{\rm P}(d^{[1]}=1\mid\bm{x})$ and $e^{[0]}(\bm{x})\equiv{\rm P}(d^{[0]}=1\mid\bm{x})$ under several assumptions, and thus the ATT can be consistently estimated. The assumptions are as follows.
@@
that is, for the expected value conditioned on $\bm{x}$ of the change in outcome over time without the treatment, the values for the treatment and control groups are the same.
@@
Events occurring between the time points of outcome collection before and after the treatment have the same impact on outcomes in both the treatment group and the control group.
@@
$0<{\rm P}(d^{[1]}=1\mid\bm{x})<1,$
that is, the propensity score is not 00 or $1$.
Abadie (2005) showed that using the weights $\rho(\bm{d},\bm{x})\equiv d^{[1]}/e^{[1]}(\bm{x})-d^{[0]}/e^{[0]}(\bm{x})$, then
@@
$={\rm E}[\Delta^{[1]}-\Delta^{[0]}\mid\bm{x},d^{[1]}=1]$
$={\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1]$
$=a(\bm{x})$ (3)
holds. Assumption 1 is used in the second equation. This equation means that by multiplying weights based on the propensity scores to the observed values of change, the conditional ATT can be estimated via DID. Then, from (2) and (3), it is shown that the optimal solution of $\bm{\theta}$ can be expressed as
@@
From this equation,
${\rm E}[e^{[1]}(\bm{x})\bm{x}\{\rho(\bm{d},\bm{x})\Delta-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$ $={\rm E}[e^{[1]}(\bm{x})\bm{x}\{{\rm E}[\rho(\bm{d},\bm{x})\Delta\mid\bm{x}]-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$
$={\rm E}[e^{[1]}(\bm{x})\bm{x}\{a(\bm{x})-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$
$=\bm{0}$ (5)
is satisfied, and therefore, the equation
@@ 2.2 Semiparametric estimation
Given an i.i.d. sample of size $n$, the $i$-th sample is denoted with the subscript $i$. Abadie (2005) proposed
@@
as a natural estimator of $\bm{\theta}^{*}$ by solving (4). When the weights based on the propensity scores are multiplied to the differences of the observed values, a term corresponding to the difference in difference appears, thus this method is called the Semiparametric DID (SDID) estimator. The following regular conditions are assumed.
@@
$(\mathrm{i})$ $\bm{\theta}^{*}$ is interior of the compact set $\bm{\Theta}\subset\mathbb{R}^{r}$. $(\mathrm{ii})$ ${\rm E}[y(t)^{2}]<\infty$ for all $t$. $(\mathrm{iii})$ ${\rm E}[\bm{x}]$ is bounded and ${\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]$ is nonsingular.
From Assumption 4 and (6),
$\hat{\bm{\theta}}=\Bigg{\{}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\Bigg{\}}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\rho(\bm{d}_{i},\bm{x}_{i})\Delta_{i}$ $\stackrel{{\scriptstyle p}}{{\to}}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}\rho(\bm{d},\bm{x})\Delta]$
$={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}a(\bm{x})]$
$={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]\bm{\theta}^{*}$
$=\bm{\theta}^{*}$
is obtained, and therefore $\hat{\bm{\theta}}$ is a consistent estimator of $\bm{\theta}^{*}$. In practice, the propensity score is often unknown. In such cases, the propensity score is modeled as $e^{[1]}(\bm{x};\bm{\alpha})={\rm logit}(\bm{x}^{{\rm T}}\bm{\alpha})$, the parameter $\bm{\alpha}$ is estimated as $\hat{\bm{\alpha}}$ by the maximum likelihood estimation (MLE), and the propensity score in (7) is replaced by $e^{[1]}(\bm{x};\hat{\bm{\alpha}})$ to obtain a semiparametric DID estimator. If the propensity score model is correctly specified, i.e., it can be written as
${\rm P}(d^{[1]}=1\mid\bm{x})={\rm logit}(\bm{x}^{{\rm T}}\bm{\alpha}^{*}),$ (8)
and if $\hat{\bm{\alpha}}$ converges to $\bm{\alpha}^{*}$ with some further assumptions, then $\hat{\bm{\theta}}$ is still a consistent estimator of $\bm{\theta}^{*}$.
3 Covariate balancing for difference-in-difference approach
3.1 Estimation method and statistical properties
Propensity scores are often estimated by the MLE, but if the model for the propensity score is misspecified, the estimator of the ATT has bias. Therefore, we deal with robust methods for the model misspecification of the propensity score. Since doubly robust estimation incorporating estimators of outcome regressions has been addressed in Sant’Anna and Zhao (2020) and Ning et al. (2020a), here we construct doubly robust estimators by incorporating covariate balancing (Imai and Ratkovic 2014). Let $\bm{{\rm O}}$ be a zero matrix and $\bm{d}=(d^{[0]},d^{[1]})$, and consider the following moment conditions:

### v3

@@ Abstract
Remarkable progress has been made in difference-in-differences (DID) approaches to causal inference that estimate the average effect of a treatment on the treated (ATT). Of these, the semiparametric DID (SDID) approach incorporates a propensity score analysis into the DID setup. Supposing that the ATT is a function of covariates, we estimate it by weighting the inverse of the propensity score. In this study, as one way to make the estimation robust to the propensity score modeling, we incorporate covariate balancing. Then, by attentively constructing the moment conditions used in the covariate balancing, we show that the proposed estimator is doubly robust. In addition to the estimation, we also address model selection. In practice, covariate selection is an essential task in statistical analysis, but even in the basic setting of the SDID approach, there are no reasonable information criteria. Here, we derive a model selection criterion as an asymptotically bias-corrected estimator of risk based on the loss function used in the SDID estimation. We show that a penalty term can be derived that is considerably different from almost twice the number of parameters that often appears in AIC-type information criteria.
Keywords: Average effect of treatment on the treated, Doubly robustness, Information criterion, Propensity score analysis, Statistical asymptotic theory
@@ 1 Introduction
The difference-in-differences (DID) approach is one of the most common methods used in fields such as economics and epidemiology to assess the effect of a particular intervention or treatment. In the basic setting, there are two groups: a treatment group receiving a specific treatment and a control group receiving no treatment or other treatment. The pre- and post-treatment outcomes of each group are measured and the change in the outcome for the control group is subtracted from the change in the outcome for the treatment group to estimate the average treatment effect on the treated (ATT). Under the parallel trend assumption, an estimation in the DID approach is valid. Recently, Abadie (2005) dealt with inference under a more natural assumption that involves using information from the covariates, specifically a parallel trend assumption conditioned on the covariates. They further proposed a semiparametric DID (SDID) approach that does not necessarily require modeling of the covariates to the outcome and that estimates the ATT by weighting the inverse of the propensity score.
@@
With or without conditionality, the parallel trend assumption is key to the DID approach. This assumption is one where the mean changes over time in the potential outcomes of the treatment group and the control group would have been similar if there had been no treatment. On the other hand, especially in epidemiology, the ignorable treatment assignment assumption has usually been invoked in causal inference (e.g. Robins et al. 1994). This is the assumption that the potential outcome variable and the treatment assignment variable, which represents the received treatment, become independent conditioned on the covariates. Although neither assumption is verifiable, to evaluate the validity of the parallel trend assumption, visual inspection is useful when observations over many time points are available. Moreover, the parallel trend assumption naturally holds if it can be supposed that there is no change over time in the expectation of outcomes without treatment. Needless to say, the parallel trend assumption is considerably weaker than the ignorable treatment assignment assumption, and estimating the ATT under such a desirable assumption is the advantage that the DID approach gains in exchange for collecting pre- and post-treatment outcomes.
@@
The methodological development of the DID approach has been growing because of the above-mentioned advantages and the increasing amount of available observational study data (see Roth et al. 2023 for a comprehensive review of recent studies on the DID approach). However, for example, Abadie (2005) assumes that the model for the propensity score is correctly specified. However, the estimator will have a bias if the model is misspecified. As an extension of Abadie (2005), Sant’Anna and Zhao (2020) assumes not only a regression model for the treatment assignment variable used to obtain the propensity score, but also a regression model for the outcome variable, and proposes an estimator for the ATT derived from the two models. Their estimator is consistent if either the treatment assignment model or the outcome regression model is correctly specified, that is, the estimator is doubly robust.
@@
In this article, as with Abadie (2005), we suppose that the ATT is a function of covariates and address the semiparametric estimation problem for that function. From the estimated function, the heterogeneity of the ATT, that is, the conditional ATT can be evaluated. Two regression models as above are considered, and in order to give double robustness to the estimation, we incorporate covariate balancing (e.g. Egel et al. 2008, Imai and Ratkovic 2014 and Fan et al. 2022), a method using propensity scores such that a certain moment condition is satisfied. This method actually estimates the treatment assignment model, but only supposes an outcome model and does not estimate it. Nevertheless, the SDID estimator derived by weighting the inverse of the propensity score, which is obtained by the covariate balancing, will be consistent if either of the regression models is correctly specified. One point to note is that, in order for the SDID approach to have this favorable asymptotic property, the moment condition has to be constructed in a different way from the usual one, which is due to considering the conditional ATT. Also, the incorporation of covariate balancing leads to an estimation via weights that always balances the moments of the covariates across groups, which is also preferable from an interpretive point of view (Imai and Ratkovic 2014). We call the proposed method covariate balancing for DID (CBD).
@@
Obviously, model selection is also an essential task in using the DID approach in a series of data-analysis tasks. Specifically, by selecting the function of covariates for the ATT addressed in this paper, it becomes possible to evaluate the heterogeneity of the ATT. There have been various important studies conducted in this area. For example, in a study estimating the effect of union membership on wages, Houngbedji (2016) treated the workers’ demographics with a linear model for causal effects and explored factors affecting the causal effect on the basis of the values of the regression coefficients and its $p$-value. In other words, they addressed the issue of covariate selection. However, there is no satisfactory information criterion, a basic tool for model selection, even for the basic setting of SDID. To remedy this situation, we will derive an information criterion for a basic SDID approach and the CBD method that is an asymptotically unbiased estimator of the risk based on the loss function used in the estimation. Note that traditional information criteria such as the generalized information criterion (GIC) (Konishi and Kitagawa 1996) cannot be applied because the loss function includes weights based on the propensity score, which are random variables. Since the SDID approach places importance on being based on the conditional parallel trend assumption, it does not rely on the ignorable treatment assignment assumption when deriving the information criterion, which causes some difficulties. As a result, we will derive a penalty term that is considerably different from twice the number of parameters, which often appears in AIC-type information criteria including the GIC.
@@
In summary, the aim of this paper is to develop a new doubly robust estimation method with the advantage of covariate balancing in the SDID approach and to derive a model selection criterion for the SDID approach, for which no reasonable model selection criteria exist yet. Its organization is as follows. Section 2 describes the model and the estimation target and introduces the SDID approach that was proposed in Abadie (2005). In Section 3, we propose the CBD method and give a theorem showing its double robustness. Numerical experiments indicate that the estimation is robust against model misspecification. In Section 4, we derive a model selection criterion for when the propensity score is known as a bias-corrected estimator of the risk that is naturally defined from the loss function used in the SDID estimation. Then, we also prove a theorem that suggests that the penalty term of the model selection criterion is asymptotically valid. As a comparison, we derive a model selection criterion that naturally extends the idea of QICw in Platt et al. (2013) to the SDID approach. Furthermore, through numerical experiments, we verify the approximation accuracy of the penalty term of the proposed criterion from the perspective of risk estimation and show its superior performance to the QICw from the viewpoint of risk minimization. Section 5 summarizes the results for the CBD method in the same way as Section 4. In other words, a model selection criterion is derived from the naturally defined risk, a theorem is given to support the validity of its penalty term, and its performance is evaluated through numerical experiments. Section 6 reports the results of the estimation and model selection for the real data in LaLonde (1986). Finally, Section 7 discusses future developments of the proposed approach.
@@ 2.1 Model and Assumptions
Suppose that there are two groups, a treatment group and a control group, and that outcomes are collected before and after the possible treatment time point, respectively. We consider the counterfactual setting, where $y^{[1]}(t)$ denotes the outcome variable of the treatment group, $y^{[0]}(t)$ the outcome variable of the control group, $t=0$ before treatment and $t=1$ after treatment. In practice, only one of the outcomes of the treatment group or control group can be obtained. By preparing treatment assignment variables such that $d^{[1]}=1\ (d^{[0]}=0)$ if the sample is assigned to the treatment group and $d^{[0]}=1\ (d^{[1]}=0)$ if the sample is assigned to the control group, the observed outcome variable collected at each time point is expressed as
@@
Denoting the change in the outcome variables at two time points in the treatment group as $\Delta^{[1]}=y^{[1]}(1)-y^{[1]}(0)$ and the change in the control group as $\Delta^{[0]}=y^{[0]}(1)-y^{[0]}(0)$, the observed change is expressed as
@@
we suppose $g(\bm{x};\bm{\theta})$ $(\bm{\theta}\in\bm{\Theta}\subset\mathbb{R}^{p})$ as a working model to approximate it, and for simplicity, set the working model to be a linear one $\bm{x}^{{\rm T}}\bm{\theta}$. Then, we estimate $\bm{\theta}$ and give the ATT by taking the expectation with respect to $\bm{x}$ in the treatment group.
Let us define the optimal value of $\bm{\theta}$ as
@@
Abadie (2005) shows that $\bm{\theta}^{*}$ can be consistently estimated with the SDID approach by using propensity scores $e^{[1]}(\bm{x})\equiv{\rm P}(d^{[1]}=1\mid\bm{x})$ and $e^{[0]}(\bm{x})\equiv{\rm P}(d^{[0]}=1\mid\bm{x})$ under several assumptions, and therefore the ATT can be consistently estimated. The assumptions are as follows.
@@
which means that, under the expectation conditioned on $\bm{x}$, the change in outcomes without treatment is the same regardless of whether it is the treatment group or the control group.
@@
Events that occur between the time points of outcome collection before and after treatment have the same impact on outcomes in both the treatment group and the control group.
@@
It holds that
$0<{\rm P}(d^{[1]}=1\mid\bm{x})<1;$
that is, the propensity score is not 0 or $1$.
Abadie (2005) showed that, by using weights $\rho(\bm{d},\bm{x})\equiv d^{[1]}/e^{[1]}(\bm{x})-d^{[0]}/e^{[0]}(\bm{x})$,
@@
$={\rm E}[\Delta^{[1]}-\Delta^{[0]}\mid\bm{x},d^{[1]}=1]={\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1]=a(\bm{x})$ (3)
holds. Assumption 1 is used in the second equality. This equation means that, by multiplying the observed values of change with weights based on propensity scores, the conditional ATT can be estimated via the DID approach. Then, from (2) and (3), the optimal solution of $\bm{\theta}$ can be expressed as
@@
Moreover, from this equation,
${\rm E}[e^{[1]}(\bm{x})\bm{x}\{\rho(\bm{d},\bm{x})\Delta-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$
$={\rm E}[e^{[1]}(\bm{x})\bm{x}\{{\rm E}[\rho(\bm{d},\bm{x})\Delta\mid\bm{x}]-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]={\rm E}[e^{[1]}(\bm{x})\bm{x}\{a(\bm{x})-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]=\bm{0}$ (5)
is satisfied, and therefore,
@@ 2.2 Semiparametric estimation
Here, for an independently and identically distributed set of samples of size $n$, the $i$-th sample will be denoted by the subscript $i$. Abadie (2005) proposed
@@
as a natural estimator of $\bm{\theta}^{*}$ by solving (4). When the observed values of change are multiplied by weights based on propensity scores, a term corresponding to the difference-in-differences appears; thus, this method is called a semiparametric DID (SDID) estimator. The following regularity conditions are assumed.
@@
$(\mathrm{i})$ $\bm{\theta}^{*}$ is in the interior of the compact set $\bm{\Theta}\subset\mathbb{R}^{r}$. $(\mathrm{ii})$ ${\rm E}[y(t)^{2}]<\infty$ for all $t$. $(\mathrm{iii})$ ${\rm E}[\bm{x}]$ is bounded and ${\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]$ is non-singular.
From Assumption 4 and (6), we obtain
$\hat{\bm{\theta}}=\bigg\{\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg\}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\rho(\bm{d}_{i},\bm{x}_{i})\Delta_{i}$
$\stackrel{{\scriptstyle p}}{{\to}}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}\rho(\bm{d},\bm{x})\Delta]$
$={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}a(\bm{x})]={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]\bm{\theta}^{*}=\bm{\theta}^{*},$
[…]
${\rm P}(d^{[1]}=1\mid\bm{x})={\rm logit}(\bm{x}^{{\rm T}}\bm{\alpha}^{*}),$
and if $\hat{\bm{\alpha}}$ converges to $\bm{\alpha}^{*}$ under some further assumptions, then $\hat{\bm{\theta}}$ is still a consistent estimator of $\bm{\theta}^{*}$.
3 Covariate balancing for difference-in-differences approach
3.1 Estimation method and its statistical properties
Propensity scores are often estimated by using MLE, but if the model for the propensity scores is misspecified, the estimator of the ATT will have bias. Here, we propose a robust method that can deal with model misspecification. While a doubly robust estimation known as augmented inverse-probability-weighted estimation, which uses estimators from an outcome regression model, has been addressed in Sant’Anna and Zhao (2020) and Ning et al. (2020a), here we construct a doubly robust estimation by incorporating covariate balancing (e.g. Egel et al. 2008, Imai and Ratkovic 2014 and Fan et al. 2022). Let $\bm{{\rm O}}$ be a zero matrix and $\bm{d}=(d^{[0]},d^{[1]})$, and let us consider the following moment conditions: